Two spherical objects have masses of 2.5 x 107 kg and 2 x 104 kg. The gravitational attraction between them is 34N. How far apart are their centers? 


1
Expert's answer
2021-11-30T17:44:53-0500

"F=G\frac{Mm}{r^2}\to r=\sqrt{\frac{GMm}{F}}=\sqrt{\frac{6.67\cdot10^{-11}\cdot 2.5\cdot10^7\cdot 2\cdot10^4}{34}}=0.99\ (m)" . Answer